OpenStudy (anonymous):

solve for x... x^2+6x+8=0 and 3x^2+6x=0

OpenStudy (anonymous):

from 3x^2+6x=0 we get x = -3x^2 / 6 = -x^2 / 2 subtituting x = -x^2/2 to x^2+6x+8=0 x^2 + 6(-x^2/2) +8 = 0

OpenStudy (anonymous):

Woah.. Im using Generic Rectangles to figure this out. Lol

OpenStudy (anonymous):

But this will still work i think.

OpenStudy (anonymous):

what is generic rectangles ? because im not native english

OpenStudy (anonymous):

\[x^2+6 x+8=(x+2) (x+4)=0 \]

OpenStudy (gw2011):

I'm assuming you want to find x for each separate equation. For: x^2+6x+8=0 (x+4)(x+2)=0 x=-4 and x=-2 For: 3x^2+6x=0 3x(x+2)=0 3x=0 x=0 and x=-2 for (x+2)

OpenStudy (anonymous):

oh that's a different equation,